OpenStudy (anonymous):

simplify sin^2 x/1-cos x

OpenStudy (anonymous):

You mean: \[\frac{ \sin ^{2}(x) }{ 1-\cos(x)}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sin^2x =1- cos^x= (1-cosx) (1+cosx) 1-cos x will vanish

OpenStudy (anonymous):

Yeah, exactly (fast typer)

OpenStudy (anonymous):

\[= \frac{ (1+\cos(x))(1-\cos(x)) }{ 1-\cos(x) }\]

OpenStudy (anonymous):

yupp thank you